OpenStudy (anonymous):

Which graph is the line (y-4)=1/4(x+2)?

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

So your equation has the point (-2,4) and a slope of 1/4

TheSmartOne (thesmartone):

which lines in your answer choice have a positive slope and cross through the point (-2,4) ?

OpenStudy (anonymous):

So it has to be positive

OpenStudy (anonymous):

It's either B or C

OpenStudy (radar):

What will y be when x=0 Figure this out from your given equation.

OpenStudy (radar):

Algebra says (0,4.5)

OpenStudy (radar):

Only one line will comply.

OpenStudy (anonymous):

Sorry I had to get my food... Um so 4?

OpenStudy (radar):

No, 4.5 or 4 1/2

OpenStudy (anonymous):

I'm confused?

TheSmartOne (thesmartone):

It has to be in between 4 and 5 exactly

OpenStudy (radar):

(y-4) = 1/4 (x + 2) y - 4 = 1/4 x + 1/2 y = (1/4) x + 4 1/2 It is just algebra.

OpenStudy (anonymous):

so B

OpenStudy (radar):

Yes B as in Boo

OpenStudy (anonymous):

Lo! Thank you again:)

OpenStudy (radar):

You're welcome, good luck with your studies.

OpenStudy (anonymous):

Thanks